# Asymptotic theory of in-context learning by linear attention

### v2

@@ Acknowledgments
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}A_{\mathrm{test}}^{\top})\qquad\text{and}\qquad\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}B_{\mathrm{test}}(\Gamma^{\ast})^{\top}),$ (SI.3.1)
where $A_{\mathrm{test}}$ and $B_{\mathrm{test}}$ are the matrices defined in (SI.2.23) and (SI.2.24), respectively. In this section, we show that the two quantities in (SI.3.1) can be obtained by studying a parameterized family of extended resolvent matrices.
@@
$\operatorname{vec}(\Gamma^{\ast})=G\left(\textstyle\sum_{\mu\in[n]}y_{\mu}\operatorname{vec}({H_{\mu}})\right)/d,$ (SI.3.2)
@@
$G=\left(\textstyle\sum_{\mu\in[n]}\operatorname{vec}({H_{\mu}})\operatorname{vec}({H_{\mu}})^{\top}/d+\tau\lambda I\right)^{-1}.$ (SI.3.3)
@@
$G(\pi)=\left(\textstyle\sum_{\mu\in[n]}\operatorname{vec}({H_{\mu}})\operatorname{vec}({H_{\mu}})^{\top}/d+\pi\Pi+\tau\lambda I\right)^{-1},$ (SI.3.4)
where $\Pi\in\mathbb{R}^{(d^{2}+d)\times(d^{2}+d)}$ is a symmetric positive-semidefinite matrix and $\pi$ is a nonnegative scalar. The original resolvent $G$ in (SI.3.3) is a special case, corresponding to $\pi=0$.
The objects in (SI.3.2) and (SI.3.4) are the submatrices of an extended resolvent matrix, which we construct as follows. For each $\mu\in[n]$, let
$z_{\mu}=\begin{bmatrix}y_{\mu}/d\\ \operatorname{vec}(H_{\mu})/\sqrt{d}\end{bmatrix}$ (SI.3.5)
@@
$\Pi_{e}=\begin{bmatrix}0&\\ &\Pi\end{bmatrix},$ (SI.3.6)
where $\Pi$ is the $(d^{2}+d)\times(d^{2}+d)$ matrix in (SI.3.4). Define an extended resolvent matrix
$G_{e}(\pi)=\frac{1}{\sum_{\mu\in[n]}z_{\mu}z_{\mu}^{\top}+\pi\Pi_{e}+\tau\lambda I}.$ (SI.3.7)
@@
$G_{e}(\pi)=\begin{bmatrix}c(\pi)&-c(\pi)q^{\top}(\pi)\\ -c(\pi)q(\pi)&G(\pi)+c(\pi)q(\pi)q^{\top}(\pi)\end{bmatrix},$ (SI.3.8)
@@
$q(\pi)\coloneqq\frac{1}{d^{3/2}}G(\pi)\left(\textstyle\sum_{\mu\in[n]}y_{\mu}\operatorname{vec}({H_{\mu}})\right)$ (SI.3.9)
@@
$\frac{1}{c(\pi)}=\frac{1}{d^{2}}\sum_{\mu\in[n]}y_{\mu}^{2}+\tau\lambda-\frac{1}{d^{3}}\sum_{\mu,\nu\in[n]}y_{\mu}y_{\nu}\operatorname{vec}(H_{\mu})^{\top}G(\pi)\operatorname{vec}(H_{\nu}).$ (SI.3.10)
By comparing (SI.3.9) with (SI.3.2), we see that
$\operatorname{vec}(\Gamma^{\ast})=\sqrt{d}\,q(0).$ (SI.3.11)
Moreover, as shown in the following lemma, the two key quantities in (SI.3.1) can also be obtained from the extended resolvent $G_{e}(\pi)$.
@@
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}A^{\top})=\frac{-1}{c(0)\sqrt{d}}\begin{bmatrix}0&\operatorname{vec}(A)^{T}\end{bmatrix}G_{e}(0)e_{1},$ (SI.3.12)
@@
$\Pi=I_{d}\otimes B$ (SI.3.13)
in (SI.3.6), then
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}B(\Gamma^{\ast})^{\top})=\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}\left(\frac{1}{c(\pi)}\right)\biggr{|}_{\pi=0}.$ (SI.3.14)
@@
The identity (SI.3.12) follows immediately from the block form of $G_{e}(\pi)$ in (SI.3.8) and the observation in (SI.3.11). To show (SI.3.14), we take the derivative of $1/c(\pi)$ with respect to $\pi$. From (SI.3.10), and using the identity
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}G(\pi)=-G(\pi)\Pi G(\pi),$ (SI.3.15)
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}\left(\frac{1}{c(\pi)}\right)$ $=\frac{1}{d^{3}}\sum_{\mu,\nu\in[n]}y_{\mu}y_{\nu}\operatorname{vec}(H_{\mu})^{\top}G(\pi)\Pi G(\pi)\operatorname{vec}(H_{\nu})$ (SI.3.16)
$=q^{\top}(\pi)\Pi q(\pi).$ (SI.3.17)
Thus, by (SI.3.11),
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}\left(\frac{1}{c(\pi)}\right)\biggr{|}_{\pi=0}$ $=\frac{1}{d}\left(\operatorname{vec}(\Gamma^{\ast})\right)^{\top}\Pi\operatorname{vec}(\Gamma^{\ast})$ (SI.3.18)
$=\frac{1}{d}\left(\operatorname{vec}(\Gamma^{\ast})\right)^{\top}\left(I_{d}\otimes B\right)\operatorname{vec}(\Gamma^{\ast}).$ (SI.3.19)
Applying the identity in (SI.1.1) to the right-hand side of the above equation, we reach (SI.3.14). ∎
@@
In light of (SI.3.13) and (SI.3.14), we will always choose
$\Pi=I_{d}\otimes B_{\mathrm{test}},$ (SI.3.20)
where $B_{\mathrm{test}}$ is the matrix defined in (SI.2.24).
SI.4 An Asymptotic Equivalent of the Extended Resolvent Matrix
In this section, we derive an asymptotic equivalent of the extended resolvent $G_{e}$ defined in (SI.3.7). From this equivalent version, we can then obtain the asymptotic limits of the right-hand sides of (SI.3.12) and (SI.3.14). Our analysis relies on non-rigorous but technically sound heuristic arguments from random matrix theory. Therefore, we refer to our theoretical predictions as results rather than propositions.
@@
$b_{\mathrm{tr}}\coloneqq\frac{1}{k}\sum_{i\in[k]}w_{i}\quad\text{and}\quad R_{\mathrm{tr}}\coloneqq\frac{1}{k}\sum_{i\in[k]}w_{i}w_{i}^{\top}$ (SI.4.1)
@@
$A_{\mathrm{tr}}\coloneqq\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}.$ (SI.4.2)
@@
$E_{\mathrm{tr}}\coloneqq\begin{bmatrix}\frac{(1+\rho)}{\alpha}I_{d}+R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\\ (1+\rho)b_{\mathrm{tr}}^{\top}&(1+\rho)^{2}\end{bmatrix}.$ (SI.4.3)
@@
Consider the extended resolvent $G_{e}(\pi)$ in (SI.3.7), with $\Pi_{e}$ chosen in the forms of (SI.3.6) and (SI.3.20). Let $\widetilde{G}_{e}$ be another matrix of the same size as $G_{e}(\pi)$. We say that $\widetilde{G}_{e}$ and $G_{e}(\pi)$ are asymptotically equivalent, if the following conditions hold.
@@
    $u^{\top}G_{e}(\pi)v\simeq u^{\top}\widetilde{G}_{e}v,$ (SI.4.4)
    where $\simeq$ is the asymptotic equivalent notation defined in (SI.1.3).
@@
    $\frac{1}{\sqrt{d}}\begin{bmatrix}0&\operatorname{vec}(A_{\mathrm{tr}})^{\top}\end{bmatrix}G_{e}(\pi)v\simeq\frac{1}{\sqrt{d}}\begin{bmatrix}0&\operatorname{vec}(A_{\mathrm{tr}})^{\top}\end{bmatrix}\widetilde{G}_{e}v.$ (SI.4.5)
@@
    Recall the notation introduced in (SI.1.2). We have
    $\frac{1}{d^{2}}\operatorname{tr}\left(\left[G_{e}(\pi)\right]_{\setminus 0}\cdot\left[I\otimes E_{\mathrm{tr}}\right]\right)=\frac{1}{d^{2}}\operatorname{tr}\left(\left[\widetilde{G}_{e}\right]_{\setminus 0}\cdot\left[I\otimes E_{\mathrm{tr}}\right]\right)+\mathcal{O}_{\prec}(d^{-1/2}),$ (SI.4.6)
@@
$\chi_{\pi}=\frac{1}{d}\operatorname{tr}\left[\Big{(}\frac{\tau}{1+\chi_{\pi}}E_{\mathrm{tr}}+\pi B_{\mathrm{test}}+\lambda\tau I_{d}\Big{)}^{-1}E_{\mathrm{tr}}\right],$ (SI.4.7)
where $B_{\mathrm{test}}$ is the positive-semidefinite matrix in (SI.2.24), with $b_{\mathrm{test}},R_{\mathrm{test}}$ chosen accroding to (SI.2.16) or (SI.2.17). The extended resolvent $G_{e}(\pi)$ in (SI.3.7) is asymptotically equivalent to
$\mathcal{G}_{e}(\pi)\coloneqq\left(\frac{\tau}{1+\chi_{\pi}}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)^{\top}\\ \frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)&I_{d}\otimes E_{\mathrm{tr}}\end{bmatrix}+\pi\Pi_{e}+\tau\lambda I\right)^{-1},$ (SI.4.8)
in the sense of Definition 1. In the above expression, $\Pi_{e}$ is the matrix in (SI.3.6) with $\Pi=I_{d}\otimes B_{\mathrm{test}}$.
In what follows, we present the steps in reaching the asymptotic equivalent $\mathcal{G}_{e}(\pi)$ given in (SI.4.8). To start, let $G_{e}^{[\mu]}$ to denote a “leave-one-out” version of $G_{e}$, defined as
$G_{e}^{[\mu]}=\frac{1}{\sum_{\nu\neq\mu}z_{\nu}z_{\nu}^{\top}+\pi\Pi_{e}+\tau\lambda I}.$ (SI.4.9)
By (SI.3.7), we have
$G_{e}\left(\textstyle\sum_{\mu\in[n]}z_{\mu}z_{\mu}^{\top}+\pi\Pi_{e}+\tau\lambda I\right)=I.$ (SI.4.10)
@@
$\chi^{\mu}(w_{\mu})\coloneqq\frac{1}{d^{2}}\operatorname{tr}\left(\left[G_{e}^{\mu}\right]_{\setminus 0}\cdot\left[I\otimes E(w_{\mu})\right]\right),$ (SI.4.13)
@@
$E(w)\coloneqq\begin{bmatrix}\frac{1+\rho}{\alpha}I_{d}+ww^{\top}&(1+\rho)w\\ (1+\rho)w^{\top}&(1+\rho)^{2}\end{bmatrix}.$ (SI.4.14)
Substituting $z_{\mu}^{\top}G_{e}^{[\mu]}z_{\mu}$ in (SI.4.11) by $\chi^{\mu}(w_{\mu})$, we get
$\sum_{\mu\in[n]}\frac{1}{1+\chi^{\mu}(w_{\mu})}G_{e}^{[\mu]}z_{\mu}z_{\mu}^{\top}+G_{e}(\pi\Pi_{e}+\tau\lambda I)=I+\Delta_{1},$ (SI.4.15)
@@
$\Delta_{1}\coloneqq\sum_{\mu\in[n]}\frac{z_{\mu}^{\top}G_{e}^{[\mu]}z_{\mu}-\chi^{\mu}(w_{\mu})}{(1+\chi^{\mu}(w_{\mu}))(1+z_{\mu}^{\top}G_{e}^{[\mu]}z_{\mu})}G_{e}^{[\mu]}z_{\mu}z_{\mu}^{\top}$ (SI.4.16)
@@
Next, we replace $z_{\mu}z_{\mu}^{\top}$ on the left-hand side of (SI.4.15) by its conditional expectation $\mathbb{E}_{w_{\mu}}\left[z_{\mu}z_{\mu}^{\top}\right]$, conditioned on the task vector $w_{\mu}$. This allows us to rewrite (SI.4.15) as
$\sum_{\mu\in[n]}\frac{1}{1+\chi^{\mu}(w_{\mu})}G_{e}^{[\mu]}\mathbb{E}_{w_{\mu}}\left[z_{\mu}z_{\mu}^{\top}\right]+G_{e}(\pi\Pi_{e}+\tau\lambda I)=I+\Delta_{1}+\Delta_{2},$ (SI.4.17)
@@
$\Delta_{2}\coloneqq\sum_{\mu\in[n]}\frac{1}{1+\chi^{\mu}(w_{\mu})}G_{e}^{[\mu]}\left(\mathbb{E}_{w_{\mu}}\left[z_{\mu}z_{\mu}^{\top}\right]-z_{\mu}z_{\mu}^{\top}\right)$ (SI.4.18)
captures the corresponding approximation error. Recall the definition of $z_{\mu}$ in (SI.3.5). Using the moment estimates in Lemma 1, we have
$\mathbb{E}_{w_{\mu}}\left[z_{\mu}z_{\mu}^{\top}\right]=\frac{1}{d^{2}}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}w_{\mu}^{\top}\otimes\begin{bmatrix}w_{\mu}^{\top}&1+\rho\end{bmatrix}\\ \frac{1}{\sqrt{d}}w_{\mu}\otimes\begin{bmatrix}w_{\mu}\\ 1+\rho\end{bmatrix}&I_{d}\otimes E(w_{\mu})\end{bmatrix}+\frac{1}{d^{2}}\begin{bmatrix}0&\\ &I_{d}\otimes\mathcal{E}_{\mu}\end{bmatrix},$ (SI.4.19)
where $E(w_{\mu})$ is the matrix defined in (SI.4.14) and
$\mathcal{E}_{\mu}=\frac{1}{\ell}\begin{bmatrix}w_{\mu}w_{\mu}^{\top}&2(1+\rho)w_{\mu}\\ 2(1+\rho)w_{\mu}^{\top}&2(1+\rho)^{2}\end{bmatrix}.$ (SI.4.20)
Replacing the conditional expectation $\mathbb{E}_{w_{\mu}}\left[z_{\mu}z_{\mu}^{\top}\right]$ in (SI.4.17) by the main (i.e. the first) term on the right-hand side of (SI.4.19), we can transform (SI.4.17) to
$\frac{\tau}{n}\sum_{\mu\in[n]}\frac{1}{1+\chi^{\mu}(w_{\mu})}G_{e}^{[\mu]}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}w_{\mu}^{\top}\otimes\begin{bmatrix}w_{\mu}^{\top}&1+\rho\end{bmatrix}\\ \frac{1}{\sqrt{d}}w_{\mu}\otimes\begin{bmatrix}w_{\mu}\\ 1+\rho\end{bmatrix}&I_{d}\otimes E(w_{\mu})\end{bmatrix}+G_{e}(\pi\Pi_{e}+\tau\lambda I)=I+\Delta_{1}+\Delta_{2}+\Delta_{3},$ (SI.4.21)

### v3

@@ Appendix C Analysis of Ridge Regression: Extended Resolvent Matrices
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}A_{\mathrm{test}}^{\top})\qquad\text{and}\qquad\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}B_{\mathrm{test}}(\Gamma^{\ast})^{\top}),$ (80)
where $A_{\mathrm{test}}$ and $B_{\mathrm{test}}$ are the matrices defined in (70) and (71), respectively. In this section, we show that the two quantities in (80) can be obtained by studying a parameterized family of extended resolvent matrices.
@@
$\operatorname{vec}(\Gamma^{\ast})=G\left(\textstyle\sum_{\mu\in[n]}y_{\mu}\operatorname{vec}({H_{\mu}})\right)/d,$ (81)
@@
$G=\left(\textstyle\sum_{\mu\in[n]}\operatorname{vec}({H_{\mu}})\operatorname{vec}({H_{\mu}})^{\top}/d+\tau\lambda I\right)^{-1}.$ (82)
@@
$G(\pi)=\left(\textstyle\sum_{\mu\in[n]}\operatorname{vec}({H_{\mu}})\operatorname{vec}({H_{\mu}})^{\top}/d+\pi\Pi+\tau\lambda I\right)^{-1},$ (83)
where $\Pi\in\mathbb{R}^{(d^{2}+d)\times(d^{2}+d)}$ is a symmetric positive-semidefinite matrix and $\pi$ is a nonnegative scalar. The original resolvent $G$ in (82) is a special case, corresponding to $\pi=0$.
The objects in (81) and (83) are the submatrices of an extended resolvent matrix, which we construct as follows. For each $\mu\in[n]$, let
$z_{\mu}=\begin{bmatrix}y_{\mu}/d\\ \operatorname{vec}(H_{\mu})/\sqrt{d}\end{bmatrix}$ (84)
@@
$\Pi_{e}=\begin{bmatrix}0&\\ &\Pi\end{bmatrix}$, (85)
where $\Pi$ is the $(d^{2}+d)\times(d^{2}+d)$ matrix in (83). Define an extended resolvent matrix
$G_{e}(\pi)=\frac{1}{\sum_{\mu\in[n]}z_{\mu}z_{\mu}^{\top}+\pi\Pi_{e}+\tau\lambda I}.$ (86)
@@
$G_{e}(\pi)=\begin{bmatrix}c(\pi)&-c(\pi)q^{\top}(\pi)\\ -c(\pi)q(\pi)&G(\pi)+c(\pi)q(\pi)q^{\top}(\pi)\end{bmatrix}$, (87)
@@
$q(\pi)\coloneqq\frac{1}{d^{3/2}}G(\pi)\left(\textstyle\sum_{\mu\in[n]}y_{\mu}\operatorname{vec}({H_{\mu}})\right)$ (88)
@@
$\frac{1}{c(\pi)}=\frac{1}{d^{2}}\sum_{\mu\in[n]}y_{\mu}^{2}+\tau\lambda-\frac{1}{d^{3}}\sum_{\mu,\nu\in[n]}y_{\mu}y_{\nu}\operatorname{vec}(H_{\mu})^{\top}G(\pi)\operatorname{vec}(H_{\nu}).$ (89)
By comparing (88) with (81), we see that
$\operatorname{vec}(\Gamma^{\ast})=\sqrt{d}\,q(0).$ (90)
Moreover, as shown in the following lemma, the two key quantities in (80) can also be obtained from the extended resolvent $G_{e}(\pi)$.
@@
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}A^{\top})=\frac{-1}{c(0)\sqrt{d}}\begin{bmatrix}0&\operatorname{vec}(A)^{T}\end{bmatrix}G_{e}(0)e_{1},$ (91)
@@
$\Pi=I_{d}\otimes B$ (92)
in (85), then
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}B(\Gamma^{\ast})^{\top})=\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}\left(\frac{1}{c(\pi)}\right)\biggr|_{\pi=0}.$ (93)
@@
The identity (91) follows immediately from the block form of $G_{e}(\pi)$ in (87) and the observation in (90). To show (93), we take the derivative of $1/c(\pi)$ with respect to $\pi$. From (89), and using the identity
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}G(\pi)=-G(\pi)\Pi G(\pi),$ (94)
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}\left(\frac{1}{c(\pi)}\right)$ $=\frac{1}{d^{3}}\sum_{\mu,\nu\in[n]}y_{\mu}y_{\nu}\operatorname{vec}(H_{\mu})^{\top}G(\pi)\Pi G(\pi)\operatorname{vec}(H_{\nu})$ (95)
$=q^{\top}(\pi)\Pi q(\pi).$ (96)
Thus, by (90),
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}\left(\frac{1}{c(\pi)}\right)\biggr|_{\pi=0}$ $=\frac{1}{d}\left(\operatorname{vec}(\Gamma^{\ast})\right)^{\top}\Pi\operatorname{vec}(\Gamma^{\ast})$ (97)
$=\frac{1}{d}\left(\operatorname{vec}(\Gamma^{\ast})\right)^{\top}\left(I_{d}\otimes B\right)\operatorname{vec}(\Gamma^{\ast}).$ (98)
Applying the identity in (39) to the right-hand side of the above equation, we reach (93). ∎
@@
In light of (92) and (93), we will always choose
$\Pi=I_{d}\otimes B_{\mathrm{test}},$ (99)
where $B_{\mathrm{test}}$ is the matrix defined in (71).
Appendix D An Asymptotic Equivalent of the Extended Resolvent Matrix
In this section, we derive an asymptotic equivalent of the extended resolvent $G_{e}$ defined in (86). From this equivalent version, we can then obtain the asymptotic limits of the right-hand sides of (91) and (93). Our analysis relies on non-rigorous but technically sound heuristic arguments from random matrix theory. Therefore, we refer to our theoretical predictions as results rather than propositions.
@@ Appendix D An Asymptotic Equivalent of the Extended Resolvent Matrix
$b_{\mathrm{tr}}\coloneqq\frac{1}{k}\sum_{i\in[k]}w_{i}\quad\text{and}\quad R_{\mathrm{tr}}\coloneqq\frac{1}{k}\sum_{i\in[k]}w_{i}w_{i}^{\top}$ (100)
@@
$A_{\mathrm{tr}}\coloneqq\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}$. (101)
@@
$E_{\mathrm{tr}}\coloneqq\begin{bmatrix}\frac{(1+\rho)}{\alpha}I_{d}+R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\\ (1+\rho)b_{\mathrm{tr}}^{\top}&(1+\rho)^{2}\end{bmatrix}$. (102)
@@
Consider the extended resolvent $G_{e}(\pi)$ in (86), with $\Pi_{e}$ chosen in the forms of (85) and (99). Let $\widetilde{G}_{e}$ be another matrix of the same size as $G_{e}(\pi)$. We say that $\widetilde{G}_{e}$ and $G_{e}(\pi)$ are asymptotically equivalent, if the following conditions hold.
@@
    $u^{\top}G_{e}(\pi)v\simeq u^{\top}\widetilde{G}_{e}v,$ (103)
    where $\simeq$ is the asymptotic equivalent notation defined in (43).
@@
    $\frac{1}{\sqrt{d}}\begin{bmatrix}0&\operatorname{vec}(A_{\mathrm{tr}})^{\top}\end{bmatrix}G_{e}(\pi)v\simeq\frac{1}{\sqrt{d}}\begin{bmatrix}0&\operatorname{vec}(A_{\mathrm{tr}})^{\top}\end{bmatrix}\widetilde{G}_{e}v.$ (104)
@@
    Recall the notation introduced in (40). We have
    $\frac{1}{d^{2}}\operatorname{tr}\left(\left[G_{e}(\pi)\right]_{\setminus 0}\cdot\left[I\otimes E_{\mathrm{tr}}\right]\right)=\frac{1}{d^{2}}\operatorname{tr}\left(\left[\widetilde{G}_{e}\right]_{\setminus 0}\cdot\left[I\otimes E_{\mathrm{tr}}\right]\right)+\mathcal{O}_{\prec}(d^{-1/2}),$ (105)
@@
$\chi_{\pi}=\frac{1}{d}\operatorname{tr}\left[\Big(\frac{\tau}{1+\chi_{\pi}}E_{\mathrm{tr}}+\pi B_{\mathrm{test}}+\lambda\tau I_{d}\Big)^{-1}E_{\mathrm{tr}}\right],$ (106)
where $B_{\mathrm{test}}$ is the positive-semidefinite matrix in (71), with $b_{\mathrm{test}},R_{\mathrm{test}}$ chosen accroding to (62) or (63). The extended resolvent $G_{e}(\pi)$ in (86) is asymptotically equivalent to
$\mathcal{G}_{e}(\pi)\coloneqq\left(\frac{\tau}{1+\chi_{\pi}}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)^{\top}\\ \frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)&I_{d}\otimes E_{\mathrm{tr}}\end{bmatrix}+\pi\Pi_{e}+\tau\lambda I\right)^{-1},$ (107)
in the sense of Definition 1. In the above expression, $\Pi_{e}$ is the matrix in (85) with $\Pi=I_{d}\otimes B_{\mathrm{test}}$.
In what follows, we present the steps in reaching the asymptotic equivalent $\mathcal{G}_{e}(\pi)$ given in (107). To start, let $G_{e}^{[\mu]}$ to denote a “leave-one-out” version of $G_{e}$, defined as
$G_{e}^{[\mu]}=\frac{1}{\sum_{\nu\neq\mu}z_{\nu}z_{\nu}^{\top}+\pi\Pi_{e}+\tau\lambda I}.$ (108)
By (86), we have
$G_{e}\left(\textstyle\sum_{\mu\in[n]}z_{\mu}z_{\mu}^{\top}+\pi\Pi_{e}+\tau\lambda I\right)=I.$ (109)
@@
$\chi^{\mu}(w_{\mu})\coloneqq\frac{1}{d^{2}}\operatorname{tr}\left(\left[G_{e}^{\mu}\right]_{\setminus 0}\cdot\left[I\otimes E(w_{\mu})\right]\right),$ (112)
@@
$E(w)\coloneqq\begin{bmatrix}\frac{1+\rho}{\alpha}I_{d}+ww^{\top}&(1+\rho)w\\ (1+\rho)w^{\top}&(1+\rho)^{2}\end{bmatrix}$. (113)
Substituting $z_{\mu}^{\top}G_{e}^{[\mu]}z_{\mu}$ in (110) by $\chi^{\mu}(w_{\mu})$, we get
$\sum_{\mu\in[n]}\frac{1}{1+\chi^{\mu}(w_{\mu})}G_{e}^{[\mu]}z_{\mu}z_{\mu}^{\top}+G_{e}(\pi\Pi_{e}+\tau\lambda I)=I+\Delta_{1},$ (114)
@@
$\Delta_{1}\coloneqq\sum_{\mu\in[n]}\frac{z_{\mu}^{\top}G_{e}^{[\mu]}z_{\mu}-\chi^{\mu}(w_{\mu})}{(1+\chi^{\mu}(w_{\mu}))(1+z_{\mu}^{\top}G_{e}^{[\mu]}z_{\mu})}G_{e}^{[\mu]}z_{\mu}z_{\mu}^{\top}$ (115)
@@
Next, we replace $z_{\mu}z_{\mu}^{\top}$ on the left-hand side of (114) by its conditional expectation $\mathbb{E}_{w_{\mu}}\left[z_{\mu}z_{\mu}^{\top}\right]$, conditioned on the task vector $w_{\mu}$. This allows us to rewrite (114) as
$\sum_{\mu\in[n]}\frac{1}{1+\chi^{\mu}(w_{\mu})}G_{e}^{[\mu]}\mathbb{E}_{w_{\mu}}\left[z_{\mu}z_{\mu}^{\top}\right]+G_{e}(\pi\Pi_{e}+\tau\lambda I)=I+\Delta_{1}+\Delta_{2},$ (116)
@@
$\Delta_{2}\coloneqq\sum_{\mu\in[n]}\frac{1}{1+\chi^{\mu}(w_{\mu})}G_{e}^{[\mu]}\left(\mathbb{E}_{w_{\mu}}\left[z_{\mu}z_{\mu}^{\top}\right]-z_{\mu}z_{\mu}^{\top}\right)$ (117)
captures the corresponding approximation error. Recall the definition of $z_{\mu}$ in (84). Using the moment estimates in Lemma 1, we have
$\mathbb{E}_{w_{\mu}}\left[z_{\mu}z_{\mu}^{\top}\right]=\frac{1}{d^{2}}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}w_{\mu}^{\top}\otimes\begin{bmatrix}w_{\mu}^{\top}&1+\rho\end{bmatrix}\\ \frac{1}{\sqrt{d}}w_{\mu}\otimes\begin{bmatrix}w_{\mu}\\ 1+\rho\end{bmatrix}&I_{d}\otimes E(w_{\mu})\end{bmatrix}+\frac{1}{d^{2}}\begin{bmatrix}0&\\ &I_{d}\otimes\mathcal{E}_{\mu}\end{bmatrix}$, (118)
where $E(w_{\mu})$ is the matrix defined in (113) and
$\mathcal{E}_{\mu}=\frac{1}{\ell}\begin{bmatrix}w_{\mu}w_{\mu}^{\top}&2(1+\rho)w_{\mu}\\ 2(1+\rho)w_{\mu}^{\top}&2(1+\rho)^{2}\end{bmatrix}$. (119)
Replacing the conditional expectation $\mathbb{E}_{w_{\mu}}\left[z_{\mu}z_{\mu}^{\top}\right]$ in (116) by the main (i.e. the first) term on the right-hand side of (118), we can transform (116) to
$\frac{\tau}{n}\sum_{\mu\in[n]}\frac{1}{1+\chi^{\mu}(w_{\mu})}G_{e}^{[\mu]}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}w_{\mu}^{\top}\otimes\begin{bmatrix}w_{\mu}^{\top}&1+\rho\end{bmatrix}\\ \frac{1}{\sqrt{d}}w_{\mu}\otimes\begin{bmatrix}w_{\mu}\\ 1+\rho\end{bmatrix}&I_{d}\otimes E(w_{\mu})\end{bmatrix}+G_{e}(\pi\Pi_{e}+\tau\lambda I)=I+\Delta_{1}+\Delta_{2}+\Delta_{3},$ (120)